An airplane makes a trip from city L to city P, which is located 213 km away from city L and on the course 148⁰. He then flies 287 km on a 75 kurs course to city F. Finally, the plane returns to city L. a. Draw the LPF triangle exactly. b. Find the distance and course from city F to city L.
1
Expert's answer
2021-09-17T00:15:49-0400

a)

triangle LPF



b)

Calculates LF, ∠F, and ∠L based on given LP, PF, and ∠P

"LF = \u221a(LP^2 + PF^2 - 2LP.PF\u00b7cos(P) )"

"⇒LF = √(213^2 + 287^2 - 2·213·287·cos(75) )\\



⇒LF = 309.99075"

Round of to nearest tenth

"\u21d2LF = 310 km"

 "∠L=arc\space cos (\frac{LF^

2

+LP

^2

−PF

^2}{

2LF·LP}


)\\

⇒ ∠L = arc\space cos (\frac{310

^2

+213

^2

−287

^2
}
{2×310×213}


)\\



⇒∠L=arc\space cos(\frac{59100}{

132060}


)\\


⇒∠L= arc\space cos(0.4475)
\\


=

∠L=63.415°"

 

Now 

"∠F = 180° − ∠L−∠P\\

⇒∠F = 180° − 63.415−75\\

⇒∠F=41.585°"


Hence Distance between L and F is 310 km

and angle from L to F 

"\u21d2360 -148-63.415-90 = 58.585\u00b0"

Now angle from F to L "= 90 - 58.585\u00b0 = 31.415\u00b0"


Hence the distance and course from city F to city L is 310 km and 31.415°